### v1

@@ 5 Simulation Example
We now present an example of a nonlinear system demonstrating the accuracy of the bound (35). In particular, we show that this bound is useful in characterizing the maximal additional energy used by a system (5) to achieve a target state, while compensating for the effect of uncontrolled inputs $u_{uc}$.
We consider the following system which contains a number of representative nonlinear functions that commonly occur in practical systems:
$\begin{split}\dot{x}_{1}&=x_{2}+\sin(x_{1})+(1+\sin^{2}(x_{2}))u_{1}+3u_{2},\\ \dot{x}_{2}&=x_{1}-x_{2}-\sin(2x_{2})+2u_{1}+(1-2\cos(x_{1}))u_{2}.\end{split}$ (36)
This system can be written in the form of equation (1), with
$f(x)=\begin{bmatrix}x_{2}+\sin(x_{1})\\ x_{1}-x_{2}-\sin(2x_{2})\end{bmatrix};\hskip 10.00002ptg(x)=\begin{bmatrix}(1+\sin^{2}(x_{2}))&3\\ 2&(1-2\cos(x_{1}))\end{bmatrix}.$
We now find constants $D_{f}$ and $D_{g}$ such that (2) and (3) are satisfied. First, $f(x)$ can be written as
$f(x)=\begin{bmatrix}0&\phantom{-}1\\ 1&-1\end{bmatrix}x+\begin{bmatrix}\sin(x_{1})\\ -\sin(2x_{2})\end{bmatrix}.$ (37)
Note that the Lipschitz constant of $f(x)$ in the $\infty$-norm satisfies (2) for the constant $D_{f}$. The Lipschitz constant of the linear term in (37) in the $\infty$-norm is $\left\|\begin{bmatrix}0&\phantom{-}1\\ 1&-1\end{bmatrix}\right\|_{\infty}=2$. Next, the Lipschitz constant of $\sin(x_{1})$ is $1$ and the Lipschitz constant of $-\sin(2x_{2})$ is $2$. Then, the Lipschitz constant of $\begin{bmatrix}\sin(x_{1})\\ -\sin(2x_{2})\end{bmatrix}$ in the $\infty$-norm is $\max\{1,2\}=2$. Combining the linear and nonlinear terms, the Lipschitz constant of $f(x)$ above is $42+\max\{1,2\}=$ 4, from which $D_{f}=4$ satisfies (2).
@@
Figure 1: Difference between nominal and malfunctioning energies, compared to the upper bound on the metric $r_{A}(t_{f},R)$. The thinner dashed lines plot $E_{T}-E_{N}$ for a variety of uncontrolled inputs $u_{uc}$.
Next, we find the constant $D_{g}$ by noting that
$g(x)=\begin{bmatrix}1&3\\ 2&1\end{bmatrix}+\begin{bmatrix}\sin^{2}(x_{2})&0\\ 0&-2\cos(x_{1})\end{bmatrix}=g_{0}+d_{g}(x).$
It is clear that $\|d_{g}(x)\|_{\infty}\leq 2$ for all $x$, and thus $D_{g}=2$.
In this example, we assume that a malfunction causes the system to lose authority over $u_{2}$, so that $u_{c}=u_{1}$ and $u_{uc}=u_{2}$, with $g(x)$ split according to (5). We fix $t_{f}=1$ and vary the distance $R$ of the initial condition from the target state, which is chosen to be $x_{tg}=[0,0]^{T}$. For notational simplicity, we denote $E_{N}\equiv E_{N}^{*}(x_{0},x_{tg},t_{f})$ for the approximate nominal energy, $E_{T}\equiv E_{T}(x_{0},x_{tg},t_{f},u_{uc})$ for the approximate total energy, and $\overline{E}_{T}\equiv\overline{E}_{T}(x_{0},x_{tg},t_{f})$ for the approximate worst-case total energy.
Fig. 1 compares the difference between the total energies and the nominal energy, to the upper bound on the metric $r_{A}(t_{f},R)$ derived in (35). This upper bound is shown as the thick red line, while the worst-case difference in energies $\overline{E}_{T}-E_{N}$ using (20) and (32) is shown as the thick blue line. It is clear that even in the worst-case, this difference in energies is upper bounded by (35). Further, the bound is meaningfully accurate, with a relatively small error. In particular, the relative error between $\overline{E}_{T}-E_{N}$ and the bound on $r_{A}(t_{f},R)$ is consistently between just $3\%$ and $4\%$ for the range of $R$ considered.
In Fig. 1, we also plot the difference between the total and nominal energies, $E_{T}-E_{N}$ for a large number of uncontrolled inputs $u_{uc}$. This includes sinusoidal inputs, constant inputs of different amplitudes, as well as decaying exponentials and sinusoids. These are shown as the thinner dashed lines. As expected, this difference is smaller than both the worst-case difference $\overline{E}_{T}-E_{N}$ and the upper bound on the metric. Once again, the worst-case difference and the upper bound on $r_{A}(t_{f},R)$ are meaningful in characterizing the additional energy required to achieve finite-time reachability under a partial loss of control authority.
We note that for $R=5$, the upper bound on $r_{A}(t_{f},R)$ is approximately $2640$ units of energy. For the different malfunctioning inputs chosen here, $E_{T}-E_{N}$ ranges from approximately from $2240$ to $2420$ units of energy. Further, the worst-case difference in energies $\overline{E}_{T}-E_{N}$ is approximately $2540$ units. Note that the upper bound on $r_{A}(t_{f},R)$ is only $3.9\%$ larger than $\overline{E}_{T}-E_{N}$. Further, $\overline{E}_{T}-E_{N}$ is at most $13\%$ larger than the actual difference in energies $E_{T}-E_{N}$ for particular choices of the uncontrolled input $u_{uc}$. Thus, $r_{A}(t_{f},R)$ is a useful quantity to approximate the increased energy required without significant conservatism, despite our knowledge of only its upper bound.
6 Conclusions
@@
In this article, we introduce a metric to bound the maximal additional energy used by a nonlinear dynamical system in achieving a finite-time reachability task, when it loses authority over a subset of its actuators. This maximal energy is compared to the nominal energy expended when no authority is lost over the inputs. We first derive an expression satisfied by the mean of the control signal used in the nominal and malfunctioning case. We then show that the energy in this mean value can be used to approximate the energy in the control signal in both cases. Under this approximation, we derive a bound on the worst-case malfunctioning energy over all possible malfunctioning inputs, and obtain an exact expression in the case when authority is lost over only one actuator. This exact expression is used to derive a bound on the maximal additional energy when authority is lost over one actuator. A simulation example on a nonlinear dynamical system demonstrates the applicability of this metric in characterizing the increased energy required, despite the approximations used for the control energies.
Our future objectives include developing a framework for quantifying energetic resilience of nonlinear systems without using approximations for nominal and malfunctioning energies. This may be possible to accomplish through the use of techniques such as control-Lyapunov functions or considering controllers with a feedback linearization structure [35].
@@ Appendix A Proof of Proposition 1
$v$ $=-\int_{0}^{t_{f}}f(x(t))\mathrm{d}t-\int_{0}^{t_{f}}d_{g}(x(t))u(t)\mathrm{d}t$
$=-t_{f}f_{0}-\int_{0}^{t_{f}}d_{f}(x(t))\mathrm{d}t-\int_{0}^{t_{f}}d_{g}(x(t))u(t)\mathrm{d}t$
@@
$\|v\|_{\infty}\leq t_{f}\|f_{0}\|_{\infty}+\left\|\int_{0}^{t_{f}}d_{f}(x(t))\mathrm{d}t\right\|_{\infty}+\left\|\int_{0}^{t_{f}}d_{g}(x(t))u(t)\mathrm{d}t\right\|_{\infty}.$
@@
$\left\|\int_{0}^{t_{f}}d_{g}(x(t))u(t)\mathrm{d}t\right\|_{\infty}\leq\int_{0}^{t_{f}}\left\|d_{g}(x(t))u(t)\right\|_{\infty}\mathrm{d}t\leq t_{f}D_{g},$
and $\left\|\int_{0}^{t_{f}}d_{f}(x(t))\mathrm{d}t\right\|_{\infty}\leq\int_{0}^{t_{f}}\|d_{f}(x(t))\|_{\infty}\mathrm{d}t\leq D_{f}\int_{0}^{t_{f}}\|x(t)-x_{0}\|_{\infty},$
@@
$\|v\|_{\infty}\leq t_{f}(\|f_{0}\|_{\infty}+D_{g})+D_{f}\int_{0}^{t_{f}}\|x(t)-x_{0}\|_{\infty}\mathrm{d}t.$ (38)
We require a method to evaluate the third term in (38). From (13),
$x(t)-x_{0}=tf_{0}+\int_{0}^{t}d_{f}(x(\tau))\mathrm{d}\tau+t_{f}B\overline{u}+\int_{0}^{t}d_{g}(x(\tau))u(\tau)\mathrm{d}\tau.$
@@
$\|x(t)-x_{0}\|_{\infty}\leq t(\|f_{0}\|_{\infty}+\|B\|_{\infty}+D_{g})+D_{f}\int_{0}^{t}\|x(\tau)-x_{0}\|_{\infty}\mathrm{d}\tau.$ (39)
Define $Q(t)\coloneqq\int_{0}^{t}\|x(\tau)-x_{0}\|_{\infty}\mathrm{d}\tau$ and $c\coloneqq\|f_{0}\|_{\infty}+\|B\|_{\infty}+D_{g}$. Note that $Q(0)=0$. In (38), we require $Q(t_{f})$ in the third term. From the first fundamental theorem of calculus [40, Theorem 5.1], $\dot{Q}(t)=\|x(t)-x_{0}\|_{\infty}$. Then, (39) reduces to
$\dot{Q}(t)\leq ct+D_{f}Q(t),~{}Q(0)=0.$ (40)
Let $Q_{1}(t)\coloneqq ct+D_{f}Q(t),Q_{1}(0)=0$ so that $\dot{Q}(t)\leq Q_{1}(t)$. Then, $\dot{Q}_{1}(t)=c+D_{f}\dot{Q}(t)\leq c+D_{f}Q_{1}(t)$. Next, let $Q_{2}(t)\coloneqq c+D_{f}Q_{1}(t),Q_{2}(0)=c$ so that $\dot{Q}_{1}(t)\leq Q_{2}(t)$. Then, $\dot{Q}_{2}(t)=D_{f}\dot{Q}_{1}(t)\leq D_{f}Q_{2}(t)$. We thus have the differential inequality
$\dot{Q}_{2}(t)\leq D_{f}Q_{2}(t),~{}Q_{2}(0)=c.$
We now invoke Grönwall’s inequality [33, Theorem 1.2.1], using which we obtain
$Q_{2}(t)\leq Q_{2}(0)e^{D_{f}t}=ce^{D_{f}t}.$
@@
$Q_{1}(t)$ $\leq\frac{c(e^{D_{f}t}-1)}{D_{f}}$
and $\text{and }Q(t)$ $\leq\frac{c(e^{D_{f}t}-1-D_{f}t)}{D_{f}^{2}}.$
@@
$\|v\|_{\infty}\leq t_{f}(\|f_{0}\|_{\infty}+D_{g})+\frac{c}{D_{f}}(e^{t_{f}D_{f}}-1-t_{f}D_{f}),$
from which the result (16) of Proposition 1 follows using the definition of $c$. $\blacksquare$

### v3

@@ 6. Simulation Examples
We now present three examples demonstrating the accuracy of the bounds (23) and (48) on the energetic resilience. In all examples, we show that these bounds are useful in characterizing the maximal additional energy used by a system (5) to achieve a target state, while compensating for the effect of uncontrolled inputs $u_{uc}$.
6.1. Example 1: Linear Driftless Model
Our first example is based on the driftless model of an underwater robot, also considered in [27]:
$\begin{bmatrix}\dot{x}\\ \dot{y}\end{bmatrix}=\begin{bmatrix}2&\phantom{-}1&\phantom{-}1\\ 0.2&-1&\phantom{-}1\end{bmatrix}\begin{bmatrix}u_{1}\\ u_{2}\\ u_{3}\end{bmatrix}$, (49)
where the input matrix is $B$. Here, $x$ and $y$ are the coordinates of the robot. The engine $u_{1}$ acts primarily along the $x$-direction, with a small bias in the $y$-direction. $u_{2}$ and $u_{3}$ are engines at a $45^{\circ}$ angle to the $x$- and $y$-directions, opposing each other. The model in [20] heavily biased the engine $u_{1}$ towards the $x$-direction, and the corresponding entry in the input matrix was $10$. We modify the entry to $2$, reducing the bias in favor of a more equitable distribution of the effect of each input.
Figure 1. Illustration of energetic resilience metric as a function of distance of the initial condition from the origin for the linear driftless model of an underwater robot. The actual difference in energies $E_{T}-E_{N}$ does not vary significantly with different choices of uncontrolled inputs $u_{uc}$, and is hence not shown here.
We assume that a malfunction causes the system to lose authority over $u_{3}$ so that $B_{c}=\begin{bmatrix}2&\phantom{-}1\\ 0.2&-1\end{bmatrix}$ and $B_{uc}=\begin{bmatrix}1\\ 1\end{bmatrix}$. In what follows, for notational simplicity, we denote $E_{N}\equiv E_{N}^{*}(x_{0},x_{tg},t_{f})$, $\overline{E}_{T}\equiv\overline{E}_{T}(x_{0},x_{tg},t_{f})$ and $E_{T}\equiv E_{T}(x_{0},x_{tg},t_{f},u_{uc})$. We now compare the nominal energy $E_{N}$ (17) to the total energy $E_{T}$ defined in (10) and worst-case total energy $\overline{E}_{M}$ in (20). In particular, we demonstrate the accuracy of the energetic resilience bound (23) in quantifying this comparison.
We fix $t_{f}=10$ and vary the distance $R$ of the initial condition $x_{0}$ from the origin. Figure 1 shows the difference between the worst-case total energy $\overline{E}_{T}$ and the nominal energy $E_{N}$ as a function of the distance $R$ of the initial condition $x_{0}$ from the origin. We also plot the upper bound on $r_{A}(t_{f},R)$ from (23) as a function of $R$, and note that this is a useful upper bound characterizing the additional energy used by the malfunctioning system. In particular, the error between the difference $\overline{E}_{T}-E_{N}$ and the upper bound on $r_{A}(t_{f},R)$ is very small. For instance, when $R=10^{2}$, the relative error between these quantities is around $20\%$. Even as $R$ increases, the relative error is consistently under $35\%$ for the entire range of $R$ considered. This example demonstrates how $r_{A}(t_{f},R)$ can be used to characterize the maximal additional energy required in a driftless system as considered in (49).
We remark that the model (49) is a minor modification of the model considered in [20]. Since that model heavily biases the effect of $u_{1}$ on $x$, this metric may be relatively less useful in that case in comparison to the model in (49). The heavier weight contributes to poor conditioning of the matrices $B^{\dagger T}B^{\dagger}$ and $B_{c}^{\dagger T}B_{c}^{\dagger}$, resulting in a more conservative bound in (23) than in the example presented here.
6.2. Example 2: Linear Model
In this example, we consider the ADMIRE fighter jet model [44], a popular benchmark for control applications [45, 15]. The linearized dynamics $\dot{x}=Ax+Bu$ for the subsystem associated with control actions were established in [44], where
$A=\begin{bmatrix}-0.9967&0&0.6176\\ 0&-0.5057&0\\ -0.0939&0&-0.2127\end{bmatrix}\text{ and }~B=\begin{bmatrix}0&-4.2423&4.2423&1.4871\\ 1.6532&-1.2735&-1.2735&0.0024\\ 0&-0.2805&0.2805&-0.8823\end{bmatrix}$. (50)
The state $x=[p,q,r]^{T}$ consists of the roll, pitch and yaw rates measured in rad/s. The control input $u=[u_{c},u_{re},u_{le},u_{r}]^{T}$ consists of deflections of the canard wings, right and left elevons and the rudder, measured in radians. Note that these inputs correspond to deflections from their equilibrium values in the linearization. We illustrate the use of the energetic resilience metric by calculating approximate control energies based only on the deflections.
Under the dynamics $\dot{x}=Ax+Bu$, where $A$ and $B$ are given in (50), we have $D_{f}=\|A\|_{\infty}=1.6143$ and $D_{g}=0$. We consider a loss of control authority over the canard wing so that $B_{uc}$ consists of the first column of $B$ in (50). It was shown in [15] that the system is resilient to the loss of this control input. We fix $t_{f}=1$ and vary the distance $R$ of the initial condition $x_{0}$ from the target state $x_{tg}$, chosen to be $0~0~0]^{T}$. We remark that while we do not directly check whether achieving this target state is always possible, tools for continuous reachability analysis [46] can be used to verify this.
@@ 6.2. Example 2: Linear Model
Figure 2. Difference between nominal and malfunctioning energies, compared to the upper bound on the metric $r_{A}(t_{f},R)$, for a linear model of a fighter jet. The thinner dashed lines plot $E_{T}-E_{N}$ for a variety of uncontrolled inputs $u_{uc}$.
Fig. 2 compares the upper bound on $r_{A}(t_{f},R)$ in (48) under the linear model for this example with the worst-case difference in energies $\overline{E}_{T}-E_{N}$ computed using the approximations (32) and (39). These quantities are plotted as a function of $R$, with $\|\tilde{x}\|_{2}\leq R$. The thinner dashed lines in this figure show the difference between the approximate total and nominal energies $E_{T}-E_{N}$ for various uncontrolled inputs $u_{uc}$, computed using (32), (38) and the definition (10). We test uncontrolled inputs ranging sinusoids of varying frequencies, constants of different amplitudes and exponential decays.
It is clear that (48) bounds the difference in energies $E_{T}-E_{N}$, even under the worst-case uncontrolled input where we have the difference $\overline{E}_{T}-E_{N}$. As $R$ increases, we note that the bound on $r_{A}(t_{f},R)$ increases faster than the difference in energies, presumably because $\overline{E}_{T}$, $E_{T}$ and $E_{N}$ all increase with $R$. This behavior contributes to smaller errors between the worst-case difference $\overline{E}_{T}-E_{N}$ and the bound on $r_{A}(t_{f},R)$ when $R$ is smaller. We further note that these quantities are of the same order of magnitude for the range of $R$ considered, even when the relative error between them is large. Fig. 2 is thus an example of how the metric $r_{A}(t_{f},R)$ can be used to estimate the maximal additional energy required to achieve a target state for a linear system.
6.3. Example 3: Wind-induced Nonlinear Model
In this example, we consider the dynamics of the ADMIRE fighter jet model subjected to wind effects. The dynamic model can then be written as $\dot{x}=Ax+Bu+f_{w}(x)$, where $A$ and $B$ are given in (50) and $f_{w}(x)$ is a nonlinear function of the state, modeling the effect of wind on the fighter jet model. We assume $f_{w}(x)$ consists of sinusoidal and constant expressions in each component as modeled in [47], and that its Lipschitz constant is no more than $1$, so that $D_{f}=\|A\|_{\infty}+1=2.6143$. In this particular instance, we choose $f_{w}(x)=\frac{1}{2}\left[\sin(p)\cos^{2}(p),~-\sin(2q),~1\right]^{T}$. We note that when $f_{w}(x)$ is unknown, we can test the applicability of the metric by searching over functions with Lipschitz constants smaller than $1$. As the dynamics are still linear in the input, we have $D_{g}=0$. As before, we consider a loss of control authority over the canard.
Figure 3. Difference between nominal and malfunctioning energies, compared to the upper bound on the metric $r_{A}(t_{f},R)$, for a nonlinear wind-affected model. The thinner dashed lines plot $E_{T}-E_{N}$ for a variety of uncontrolled inputs $u_{uc}$.
In Fig. 3, we compare the upper bound on $r_{A}(t_{f},R)$ in (48) to the worst-case difference in energies $\overline{E}_{T}-E_{N}$, where we use the approximations (32) and (39) to compute these quantities. We plot these quantities as a function of the distance $R$ of the initial condition $x_{0}$ from the target state $x_{tg}=[0~0~0]^{T}$, and fix the final time as $t_{f}=1$. As before, tools for reachability analysis can be used to verify whether the target state can be achieved under these settings. The thinner dashed lines denote the difference in energies $E_{T}-E_{N}$ for different uncontrolled inputs $u_{uc}$, computed using the approximations (32), (38) and the definition (10). As before, the uncontrolled inputs we test consist of sinusoids of varying frequencies, exponential decays and constant inputs of different amplitudes. It is clear that even under the worst-case uncontrolled input, the difference in energies $\overline{E}_{T}-E_{N}$ is bounded by (48). As expected, the actual differences in energies $E_{T}-E_{N}$ for different uncontrolled inputs is also bounded by (48).
While the relative error between the worst-case difference $\overline{E}_{T}-E_{N}$ and the bound on $r_{A}(t_{f},R)$ is large, we note that these quantities are of the same order of magnitude. This error is also smaller when $x_{0}$ is closer to $x_{tg}$, i.e., when $R$ is small. In this case, the bound on $r_{A}(t_{f},R)$ is less than $3$ times larger than the worst-case difference in energies. We also note that the results for the nonlinear model in Fig. 3 show an improvement over the results for the linear model in Fig. 2. This is due to the nonlinear function $f_{w}(x)$ that models wind effects. The energy required in the control signal can decrease when this function acts in the same direction that the system seeks to move towards.
Figure 4. Sensitivity of metric to different amplitudes of wind disturbance.
In Fig. 4, we illustrate the sensitivity of the metric to different amplitudes of wind disturbances. In particular, we test disturbances of the form $f_{w}(x)=\frac{c}{2}\left[\sin(p)\cos^{2}(p),~-\sin(2q),~1\right]^{T}$, where $c$ is a varying parameter signifying the amplitude of disturbances. Note that $c=1$ is the case considered in Fig. 3. Clearly, the upper bound on the metric is sensitive to the choice of $c$, and for larger values of $c$, may exhibit more conservative behavior.
We remind the reader that the expressions plotted in Figs. 2 and 3 are based on approximations made in (32) and (38), resulting in approximations for $\overline{E}_{T}$ (39) and $r_{A}(t_{f},R)$ (48). Under these approximations, we note that $r_{A}(t_{f},R)$ is a useful quantity that estimates the increased energy required without significant conservatism, despite our knowledge of only its approximate upper bound.
7. Conclusions
@@
In this article, we introduce a metric to bound the maximal additional energy used by a nonlinear dynamical system in achieving a fixed-time reachability task, when it loses authority over a subset of its actuators. This maximal energy is compared to the nominal energy expended when no authority is lost over the inputs. We first consider the special case of linear driftless systems and recall the energies used by the control signal in both the nominal and malfunctioning cases. This analysis is used to derive a bound on the metric to quantify the maximal additional energy used in the malfunctioning case, assuming control authority is lost over one actuator. For the general nonlinear case, we first derive an expression satisfied by the mean of the control signal used in the nominal and malfunctioning case. We then show that the energy in this mean value can be used to approximate the energy in the control signal in both cases. Under this approximation, we derive a bound on the worst-case malfunctioning energy over all possible malfunctioning inputs. In the case when authority is lost over only one actuator, this bound reduces to an equality under the aforementioned approximation. This expression is finally used to derive a bound on the resilience metric when authority is lost over one actuator. A set of simulation examples on driftless, linear and nonlinear dynamical systems demonstrate the applicability of this metric in characterizing the increased energy required, despite the approximations used for the control energies in the nonlinear case.
We note that we have not numerically simulated optimal controllers for nonlinear systems, and thus do not compare the metric with the difference between the true optimal controllers for the nominal and malfunctioning systems. Such a numerical simulation is an avenue to explore in order to further demonstrate the usefulness of this metric. Another important avenue to explore is the actual design of resilient controllers. Such a resilient design must also consider that controlled and uncontrolled components may not be known in advance, and must adapt to changes in dynamics. A potential approach towards addressing this problem is presented in [48] under a switched systems framework. Combining that approach with the quantification of additional energy in this paper is a useful avenue to explore. We also note that it is useful to study the case of losing authority over multiple actuators, through simulation or theory, though we may only obtain a more conservative bound in (48).
@@ Appendix A Proof of Proposition 2
$v$ $=-\int_{0}^{t_{f}}f(x(t))\mathrm{d}t-\int_{0}^{t_{f}}d_{g}(x(t),x_{0})u(t)\mathrm{d}t=-t_{f}f_{0}-\int_{0}^{t_{f}}d_{f}(x(t),x_{0})\mathrm{d}t-\int_{0}^{t_{f}}d_{g}(x(t),x_{0})u(t)\mathrm{d}t$
@@
$\|v\|_{\infty}\leq t_{f}\|f_{0}\|_{\infty}+\left\|\int_{0}^{t_{f}}d_{f}(x(t),x_{0})\mathrm{d}t\right\|_{\infty}+\left\|\int_{0}^{t_{f}}d_{g}(x(t),x_{0})u(t)\mathrm{d}t\right\|_{\infty}.$
@@
$\left\|\int_{0}^{t_{f}}d_{g}(x(t),x_{0})u(t)\mathrm{d}t\right\|_{\infty}$ $\leq\int_{0}^{t_{f}}\left\|d_{g}(x(t),x_{0})u(t)\right\|_{\infty}\mathrm{d}t$
$\leq\int_{0}^{t_{f}}\|d_{g}(x(t),x_{0})\|_{\infty}\mathrm{d}t\leq D_{g}\int_{0}^{t_{f}}\|x(t)-x_{0}\|_{\infty}\mathrm{d}t,$
$\text{and }\left\|\int_{0}^{t_{f}}d_{f}(x(t),x_{0})\mathrm{d}t\right\|_{\infty}$ $\leq\int_{0}^{t_{f}}\|d_{f}(x(t),x_{0})\|_{\infty}\mathrm{d}t\leq D_{f}\int_{0}^{t_{f}}\|x(t)-x_{0}\|_{\infty}\mathrm{d}t,$
@@
$\|v\|_{\infty}\leq t_{f}\|f_{0}\|_{\infty}+(D_{f}+D_{g})\int_{0}^{t_{f}}\|x(t)-x_{0}\|_{\infty}\mathrm{d}t.$ (51)
We require a method to evaluate the second term in (51). From (25),
[…]
@@
[…]
$\dot{Q}(t)\leq ct+(D_{f}+D_{g})Q(t),~Q(0)=0.$ (53)
Let $Q_{1}(t)\coloneqq ct+(D_{f}+D_{g})Q(t),Q_{1}(0)=0$ so that $\dot{Q}(t)\leq Q_{1}(t)$. Then, $\dot{Q}_{1}(t)=c+(D_{f}+D_{g})\dot{Q}(t)\leq c+(D_{f}+D_{g})Q_{1}(t)$. Next, let $Q_{2}(t)\coloneqq c+(D_{f}+D_{g})Q_{1}(t),Q_{2}(0)=c$ so that $\dot{Q}_{1}(t)\leq Q_{2}(t)$. Then, $\dot{Q}_{2}(t)=(D_{f}+D_{g})\dot{Q}_{1}(t)\leq(D_{f}+D_{g})Q_{2}(t)$. We thus have the differential inequality
$\dot{Q}_{2}(t)\leq(D_{f}+D_{g})Q_{2}(t),~Q_{2}(0)=c.$
We now invoke Grönwall’s inequality [38, Theorem 1.2.1], using which we obtain
$Q_{2}(t)\leq Q_{2}(0)e^{(D_{f}+D_{g})t}=ce^{(D_{f}+D_{g})t}.$
@@
$Q_{1}(t)$ $\leq\frac{c(e^{(D_{f}+D_{g})t}-1)}{D_{f}+D_{g}}$
$\text{and }Q(t)$ $\leq\frac{c(e^{(D_{f}+D_{g})t}-1-(D_{f}+D_{g})t)}{(D_{f}+D_{g})^{2}}.$
@@
$\|v\|_{\infty}\leq t_{f}\|f_{0}\|_{\infty}+\frac{c}{D_{f}+D_{g}}\left(e^{t_{f}(D_{f}+D_{g})}-1-t_{f}(D_{f}+D_{g})\right).$
Substituting $c=\|f_{0}\|_{\infty}+\|B\|_{\infty}$ and simplifying,
$\|v\|_{\infty}\leq\overline{v}\coloneqq\frac{1}{D_{f}+D_{g}}\left[c\left(e^{t_{f}(D_{f}+D_{g})}-1\right)-t_{f}(D_{f}+D_{g})\|B\|_{\infty}\right],$
thus concluding the proof. $\blacksquare$
Appendix B Table of Non-Approximate and Approximate Expressions
Table 1 provides a classification of which equations are exact, i.e., non-approximate expressions or upper bounds, and which equations are based on the approximation from Lemma 1. We note that all expressions for linear driftless systems are non-approximate, and all expressions for nonlinear systems are approximate.
Table 1. Classification of Non-Approximate and Approximate Expressions
Dynamic Model Classification Equations
Linear Driftless System Non-Approximate Equalities (17), (18), (20)
Non-Approximate Upper Bounds (19), (23)
Nonlinear System Approximate Equalities (32), (38), (44)
Approximate Upper Bounds (39), (48)